OpenStudy (anonymous):

Hi guys I would like to know which of these volume is greater (A) 251.33 (B) 169.64 .........Ive gotten the answers by using the formula V=Pi Ratio^2 H.....but I dont which cylinder volume is greater.

Directrix (directrix):

Would you post the original problem from which you got the two volumes? >>>(A) 251.33 (B) 169.64 I think it would be good to check those before choosing an answer for the larger volume.

OpenStudy (anonymous):

number 2

Directrix (directrix):

So, you did these two calculations: V = π * r² * h V = π * 4² * 5 V = 80 π cubic inches and V = π * r² * h V = π * 3² * 6 V = 54 π cubic inches Which of these is the larger volume? @June_Berry

OpenStudy (anonymous):

80TT

Directrix (directrix):

Correct and that approximates to 251.327. Your posted volumes are correct. Good work.

Directrix (directrix):

That is the first volume and not the second one. The units of measurement are the same. Therefore, you can directly compare 251 and 169 to choose the larger volume.

OpenStudy (anonymous):

thanks appreciate it!

Directrix (directrix):

You are welcome.